Simplify (a^2-b^2)/(a^2+ab)
1
Expert's answer
2020-08-10T18:11:16-0400

(a2 - b2)/(a2 + ab)

factorize;

(a2 - b2)=(a - b)(a + b)

(a2 + ab)=a(a + b)

Therefore;

(a2 - b2)/(a2 + ab)=(a - b)(a + b)/(a(a + b)) then (a +b) vanishes leaving


=(a - b)/a